The condensed financial statements for OIL Inc. and ERS Company for the year ended December 31, Year 5, follow:

OIL ERS
Revenues $ 945,000 $ 330,000
Expenses 675,000 215,000
Net income $ 270,000 $ 115,000
Retained earnings, 1/1/Year 5 $ 815,000 $ 215,000
Net income 270,000 115,000
Dividends paid 105,000 0
Retained earnings, 12/31/Year 5 $ 980,000 $ 330,000
Cash $ 95,000 $ 125,000
Receivables and inventory 415,000 185,000
Patented technology (net) 915,000 322,500
Equipment (net) 715,000 615,000
Total assets $ 2,140,000 $ 1,247,500
Liabilities $ 615,000 $ 432,500
Common shares 545,000 485,000
Retained earnings 980,000 330,000
Total liabilities and equities $ 2,140,000 $ 1,247,500

On December 31, Year 5, after the above figures were prepared, OIL issued $255,000 in debt and 15,000 new shares to the owners of ERS for 80% of the outstanding shares of that company. OIL shares had a fair value of $35 per share.

OIL also paid $45,000 to a broker for arranging the transaction. In addition, OIL paid $47,000 in stock issuance costs. ERS’s equipment was actually worth $720,000, but its patented technology was appraised at only $295,000.

Required:

What are the consolidated balances for the year ended/at December 31, Year 5, for the following accounts?

(a) Net income

OIL's net income considered in the Consolidated Financial Statement            $

(b) Retained earnings, 1/1/Year 5

OIL's retained earnings in the Financial statement for Consolidation            $

(c) Equipment

Value of equipment after acquisition            $

(d) Patented technology

Value of patent after acquisition            $

(e) Goodwill

Goodwill            $

(f) Liabilities

Total liabilities after acquisition            $

(g) Common shares

Total value of common shares after acquisition            $

(h) Non-controlling interests

Total value of non-controlling interest after acquisition            $
Answer #1

Part A

OIL's net income considered in the Consolidated Financial Statement = Parent's income prior to the date of acquisition- paid to broker = 270000-45000 = $225000

Part B

OIL's retained earnings in the Financial statement for Consolidation = Parent's retained earnings prior to the date of acquisition, at 1/1/Year 5 = $815000

Part C

Value of equipment after acquisition = 715000+615000+105000 = $1435000

Part D

Value of patent after acquisition = 915000+322500-27500 = $1210000

Part E

Goodwill = $82500

Part F

Total liabilities after acquisition = 615000+432500+255000 = 1302500

Part G

Total value of common shares after acquisition = 545000+(15000*35)-47000= $1023000

Part H

Total value of non-controlling interest after acquisition = 20%*975000 = $195000

Cost of 80% of ERS ($255,000 + 15,000 shares x $35)

780000

Implied value of 100% of ERS

975000

Carrying amount of ERS’s net assets

Assets

1247500

Liabilities

432500

815000

160000

Acquisition differential

Allocated:

Equipment (720000-615000)

105000

Patent technology (295000-322500)

-27500

77500

Goodwill

$82500